OpenStudy (anonymous):

Given the equation y=(the square root of) x+9, what domain will produve a range that is in the set of real numbers

OpenStudy (anonymous):

Think about it another way: what domain will give us non-real (imaginary) numbers? As it's a square root, we're looking to square root a negative number so x<-9. As we're looking for real numbers, take the opposite of this and we get x>=-9 :) You can double check your domain: x=-9 => f(-9)=0 Real x=-8 => f(-8)=+-1 Real x=-7 => f(-7)=sqrt(2) Real And so on, we can clearly see the numbers are getting bigger and remaining real.

OpenStudy (anonymous):

Neither, the function gives real numbers when x is greater than or equal to -9. x>=-9

OpenStudy (anonymous):

What is the range? Your function is f(x)=(x-5)/(x+8), yes?

OpenStudy (anonymous):

yeah, the range is the set of real numbers

OpenStudy (anonymous):

and then also my last question is given the relation y=x^2-4 state the range if the domain is {-2, -1, 0, 1, 2}

OpenStudy (anonymous):

Your domain is just x≠-8 as you don't want to divide by 0. Any other number gives a real number as an output - try some values of x in the expression and you'll see. As for the other one. Put each number in: f(x)=x^2-4 f(-2)=0 f(-1)=-3 f(0)=-4 f(1)=-3 f(2)=0 So our range is {-4, -3, 0}

OpenStudy (anonymous):

thanks so much!! and sorry this is the last question but theres one that says state the range if the domain is {x-2 < or equal to x < or equal to 2}

jimthompson5910 (jim_thompson5910):

so the domain is \[\Large -2 \le x \le 2\] right?

OpenStudy (anonymous):

it says x-2 for the first one

OpenStudy (anonymous):

theres just an x before the -2

jimthompson5910 (jim_thompson5910):

ok so \[\Large x-2 \le x \le 2\] right?

OpenStudy (anonymous):

yeah

jimthompson5910 (jim_thompson5910):

ok you have to break up the compound inequality into these two inequalities \[\Large x-2 \le x\] and \[\Large x \le 2\]

jimthompson5910 (jim_thompson5910):

what do you get when you solve \[\Large x-2 \le x\] for x

OpenStudy (anonymous):

not sure...would it be -2?

jimthompson5910 (jim_thompson5910):

well if you subtract x from both sides, you get \[\Large -2 \le 0\] which is always true

OpenStudy (anonymous):

right

jimthompson5910 (jim_thompson5910):

so it turns out the solution to \[\Large x-2 \le x \le 2\] is \[\Large x \le 2\]

jimthompson5910 (jim_thompson5910):

your domain is \[\Large x \le 2\]

OpenStudy (anonymous):

thanks so much! got it!

jimthompson5910 (jim_thompson5910):

use that to find the range

OpenStudy (anonymous):

how would you do that?

OpenStudy (anonymous):

is it < or equal to 3?

jimthompson5910 (jim_thompson5910):

what is the original function again? can you draw it out?

OpenStudy (anonymous):

\[y=x ^{2}\]

jimthompson5910 (jim_thompson5910):

I thought it was a square root or something?

OpenStudy (anonymous):

oh no sorry that was a different one! i thought you were helping with the second question

OpenStudy (anonymous):

the question is state the range if the domain is \[x-2\le x \le 2\]

jimthompson5910 (jim_thompson5910):

ok so as shown above, the domain simplifies to \[\Large x \le 2\]

jimthompson5910 (jim_thompson5910):

so graph y = x^2 using a graphing calculator like this one https://www.desmos.com/calculator

jimthompson5910 (jim_thompson5910):

however, only graph when x <= 2 so you will type in "x^2{x<=2}" without quotes the {x<=2} ensures only the portion to the left of x=2 is graphed (x=2 is included)

jimthompson5910 (jim_thompson5910):

you then use the graph to determine the range